OpenStudy (anonymous):

ben is using a map to plan a hike the map is based on a scale of 1:80,000 in cm if the route is 20 cm on the map how any kilometers is the actual hike?

OpenStudy (anwaarullah):

20/80,000=2/8000 =1/4000 =.00025 km